If c is the measure of the hypothenuse, find the missing measure. Round to the nearest tenth, if necessary.
a=?
b=16
c=43

Answer by Theo(13342)

If c is the hypotenuse of a right triangle, then a and b are the legs.
The pythagorean formula states that:

c^2 = a^2 + b^2

If c = 43 and b = 16 then this formula becomes:

43^2 = a^2 + 16^2 which becomes

1849 = a^2 + 256

Subtract 256 from both sides of the equation to get

1593 = a^2
Take the square root of both sides of this equation to get:

39.91240409 = a

To confirm, plug this in the original pythagorean formula of:

c^2 = a^2 + b^2 to get:

43^2 = 39.91240409^2 + 16^2 which becomes:

1849 = 1593 + 256 which becomes

1849 = 1849 confirming that the value of a is good.